solve by writing and solving an equation
The larger of two numbers is three times the smaller. The sum of the two numbers is 6. What are the numbers?

Answer by TimothyLamb(4379)

---
y = 3x
x + y = 6
---
put the system of linear equations into standard form:
---
3x - y = 0
x + y = 6
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 1.5
y = 4.5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php